Santa Cruz Bottling is a manufacturer of organic soft drinks on the coast of central California. Its products are enjoying a growing reputation and increased demand throughout the American Southwest. Because of the high cost of transporting soft drinks, it is considering a new plant to serve the states of New Mexico and Arizona. A key concern in its search for a new location is the resultant transportation costs to serve its key markets. Following is a list of cities where its main wholesale customers are located, along with estimated annual demand in cases of product for each.

CITY x-COORDINATES y-COORDINATES NUMBER OF CASES
Phoenix 270 270 22,500
Tucson 370 145 18,000
Albuquerque 820 470 26,000
Santa Fe 870 535 14,500


a. Use the centroid method to recommend a location for the new bottling plant. (Round your answers to 1 decimal place.)

Cx=

Cy=
Answer #1

Answer: - According to given data

Formula to calculate cordinates = sum of ( cordinate*number of cases) / Total number of cases

For Cx = 270(22,500)+370(18,000)+820(26,000)+870(14,500) / 81,000 = 576.2

For Cy = 270(22,500)+145(18,000)+470(26,000)+535(14,500) / 81,000 = 353.9

So Cx = 576.2 and Cy = 353.9